OpenStudy (anonymous):

What is the 7th number in the pattern: 1, 4, 9, 16, 25, …?

OpenStudy (anonymous):

find the pattern.....

OpenStudy (anonymous):

1^2 = 1 2^2 = 4 3^2 = 9......

OpenStudy (anonymous):

think of it as 1x1=1, 2x2=4, 3x3=9,4x4=16,5x5=25,.......keep that pattern going like JuanitaM said... (just simplified)